A quantity of a fluid with an initial pressure of 500KPa and an initial volume of 1m3 is made to undergo as process w/in a piston-cylinder system, Compute the work done by the fluid according to the condition that pressure is proportional to the square of volume. Express your answer in J.


1
Expert's answer
2022-01-05T09:36:02-0500

Determine how exactly pressure depends on volume:


"P(v)=kv^2,\\\space\\
500000=k·1^2,\\\space\\
k=500·10^3"

The work is

"W=\int^{V_2}_{V_1}P\text dv=\int^{2}_{1}kv^2\text dv=\dfrac13kv^3,\\\space\\
W=500·10^3·\dfrac13(2^3-1^3)=1166667\text{ J}."